OpenStudy (anonymous):

The perimeter of a triangle is 28ft. The medium side is 4ft longer than the shortest side,while the longest side is twice as long as the shortest side. What are the lengths of the three sides?

OpenStudy (anonymous):

I came up with 2c+ (4+c) +c=28 2c+ 4+c+c=28 4c+4=28 4c=24 divide by 4 c=6 a=12 b=10 can someone tell me if this is correct??

OpenStudy (texaschic101):

P = a + b + c 28 = c + c + 4 + 2c -- commbine like terms 28 = 4c + 4 28 - 4 = 4c 24 = 4c 24/4 = c 6 = c shortest side (c) = 6 medium side = c + 4 = 6 + 4 = 10 longest side = 2c = 2(6) = 12 check.. 28 = 12 + 10 + 6 28 = 28 (correct) you got it :)